# Leveraging Online Data to Enhance Medical Knowledge in a Small Persian Language Model

### v2

@@ Abstract
The rapid advancement of language models has demonstrated the potential of artificial intelligence in the healthcare industry. However, small language models struggle with specialized domains in low-resource languages like Persian. While numerous medical-domain websites exist in Persian, no curated dataset or corpus has been available—making ours the first of its kind. This study explores the enhancement of medical knowledge in a small language model by leveraging accessible online data, including a crawled corpus from medical magazines and a dataset of real doctor-patient Q&A pairs. We fine-tuned a baseline model using our curated data to improve its medical knowledge. Benchmark evaluations demonstrate that the fine-tuned model achieves improved accuracy in medical question answering and provides better responses compared to its baseline. This work highlights the potential of leveraging open-access online data to enrich small language models in medical fields, providing a novel solution for Persian medical AI applications suitable for resource-constrained environments.
Index Terms:
persian medical question answering,small language model,medical language models, data crawling
@@ I Introduction
In response to these challenges, we present a novel approach with our model, Gaokerena 111 Our language model is named after Gaokerena, an ancient Persian mythological tree believed to possess healing properties and grant immortality to those who consume its fruit. , which fine-tunes a baseline model, aya-expanse-8b [5] , on the crawled data comprising a persian medical corpus and a medical free form farsi question answering dataset. Importantly, our model, corpus, and datasets are all open-source, promoting transparency and collaboration within the research community. This development aims to enhance access to Persian medical information and support secure, efficient interactions within the healthcare environment. By bridging the existing gaps in resources and leveraging advancements in NLP, our work contributes to the growing landscape of AI in medicine, particularly for Persian-speaking users.
@@ II-B Related Works in Persian
As previously mentioned, there has been limited research focused on Persian medical language models, highlighting a significant gap in resources for the Persian-speaking medical community. Furthermore, existing works on Persian medical question-answering systems are entirely closed-source regarding their datasets, models, and codebases. This lack of public resources leaves the field largely underexplored, presenting researchers with an almost blank slate to build upon. On the other hand, all of these efforts have primarily concentrated on extractive solutions, which aim to retrieve relevant information from predefined sources, rather than employing generative approaches capable of producing context-aware responses.
Perhaps the most notable effort in this area is Sina-BERT [4], which involved training a BERT model [11] using a crawled corpus alongside Persian annotated datasets specifically developed for various tasks, including medical question answering, medical sentiment analysis, and medical question retrieval. Sina-BERT is the most similar work to ours among Persian language focused efforts; however, it differs in that it uses a BERT model—an encoder-based language model—as its baseline. This choice limits its capability for generative AI tasks, as BERT is primarily designed for understanding and extracting information rather than generating answers.
@@ III Baseline Model
We have chosen aya-expanse-8b as our baseline model primarily due to the lack of open-source Persian medical language models, which necessitates the use of a general-purpose language model. While there are several multilingual options available, including aya-expanse [5], Gemma2 [15], Qwen2 [16], and PersianMind [17], we have determined that aya-expanse is the most suitable choice for our needs. One key reason is that the training data for the other models predominantly consists of non-Persian languages, leading to biases that may result in the generation of non-Persian characters, even when we explicitly instruct the model to use only Persian. In contrast, aya-expanse demonstrates a robust understanding of Persian grammar and produces grammatically rich Persian text, making it a more reliable option for our research. Furthermore, if we merge our updated parameters into aya-vision [18] , another model of aya family, instead of aya-expanse, we gain the capability to incorporate medical images such as MRIs and CT scans as inputs, thereby enhancing our model’s applicability to the medical domain.
@@ IV-A Corpus
I. Garcia Ferrero et al. [19] collected medical corpora dedicated to four languages (English, French, Spanish, and Italian), which can be compared to ours, as shown in Table I. The accompanying Figure 1 illustrates the share of each magazine within our corpus, effectively highlighting the diversity of sources and underscoring the need to address gaps in available resources to foster innovation and improve health-related applications.
TABLE I: Comparison of Our Corpus with Corpora Collected by I. Garcia Ferrero et al.
@@ IV-B Dataset
The collection of a real-world doctor-patient question-answering dataset is crucial for enhancing the capabilities of language models in the healthcare domain. Such a dataset allows models to learn valuable information derived from authentic interactions between healthcare providers and patients. By analyzing these real-world exchanges, language models can grasp the nuances of medical terminology, patient concerns, and the context surrounding healthcare inquiries. Furthermore, this dataset equips models with the ability to learn not just the factual content of responses but also the appropriate structure and tone for answering questions. This dual learning process is essential, as it enables the model to generate accurate, empathetic, and contextually relevant responses, ultimately improving patient communication and support in medical environments. In this context, Yang Liu [20] highlights several real-world doctor-patient question-answering datasets in his survey, a comparison of these datasets with ours can be found in Table II. In an era where technology increasingly aids healthcare, a robust doctor-patient dataset stands as a foundational element in training models that can effectively contribute to better healthcare delivery.
TABLE II: Comparison of Our dataset with others
@@
In our research, we crawled more than 180,000 question-answer pairs from Persian medical forums, employing both manual and automatic filtering methods to refine the dataset through a laborious cleaning process. This approach is similar to the work done by Yunxiang Li et al. [6]. in their article on the Chat Doctor medical language model, where they extracted data from English medical forums. Notably, Yunxiang Li discarded about half of the question-answer pairs based on the length of the answers, as shorter responses are generally inadequate for training a model and can lead the model to learn to provide brief answers. However, we faced a greater challenge; Persian doctors tend to provide much shorter answers compared to their English counterparts, resulting in the necessity to discard over 80% of our question-answer records to ensure quality and relevance for our training purposes.
As you can see in the Figure 2 to create our dataset, we utilized patient-doctor interactions from the drhast and niniban platforms for the training split. It is important to note that drhast does not provide all of its doctor-patient interaction records on its site; it only offers access to the last 2,000 records. Additionally, each record is linked to 100 related records, complicating the crawling process. To address this challenge, we treated their data as a graph and performed a breadth-first search, which took about two weeks to extract 120,000 records out of a total of 200,000. For the test set, we used the doctor-yab and isovisit sites, ensuring diversity by translating the K-QA question-answering dataset [27] and appending it to our test split. This comprehensive approach not only enriched our dataset but also underscored the importance of real-world doctor-patient interactions in training effective language models.
@@ V-A Fine tuning
We fine-tuned the 8-billion-parameter checkpoint of the aya-expanse model on 60% of our corpus, focusing on minimizing resource usage. To ensure an efficient fine-tuning process, we employed gradient checkpointing and a small batch size of 2, reducing memory requirements during training. Additionally, we used gradient accumulation steps of 16, effectively increas- ing the overall batch size to 32 and enabling stable training dynamics. To further reduce the memory usage of our fine- tuning process, we leveraged Low Rank Adaptation (LoRA) [28] to significantly reduce the number of trainable parameters. Specifically, we implemented a rank of 8, an alpha value of 16, and a dropout rate of 5%, applying the LoRA weights to all trainable parameters of each transformer layer.
To further optimize the process, we employed efficient tokenization and memory-aware training techniques. The to- kenization process divided the input text into manageable token sequences, ensuring consistent input and label structures by truncating, padding, and handling overflowed tokens to maintain contextual integrity within a fixed context length. This streamlined preparation, coupled with the LoRA-based fine-tuning, was further enhanced by Flash Attention 2 [29]. By minimizing memory overhead, Flash Attention 2 allowed us to handle longer context lengths and larger batch sizes efficiently, enabling effective fine-tuning for the next-token prediction objective while balancing computational efficiency with model performance.
@@ V-B Instruction tuning
Following the fine-tuning stage, we performed instruction- tuning on the fine-tuned model using our crawled free-form Farsi question-answering (MF3QA) dataset. This stage utilized the default template of the aya-expanse model and retained the same techniques and almost the same hyperparameters as the fine-tuning stage, with a few adjustments. Specifically, we employed the LoRA method with a rank of 2, an alpha value of 2, and a dropout rate of 0.4, and increased the weight decay to 0.5 instead of 0.1. The instruction-tuning process was conducted for a single epoch, enabling the model to better understand and generate responses tailored to Farsi question- answering tasks. This targeted optimization further refined the model’s capabilities, enhancing its effectiveness on our specific dataset.
@@ V-C Carbon Footprint
The carbon footprint of our model optimization—including both fine-tuning and instruction-tuning—was estimated based on hardware specifications and operational duration. The pro- cess ran for a combined total of 19 hours on a NVIDIA A100 PCIe 40GB GPUs hosted in Google Cloud Platform’s asia- east1 region. Assuming a typical power consumption of 250 watts per GPU, the total energy used was 4.75 KWh (250 watts × 19 hours). Using the carbon intensity factor of the asia-east1 grid (0.56 kilograms of CO2 equivalent per kWh), this translates to 2.66 kilograms of CO2 equivalent emitted during the tuning process [30].
@@ VI Results
In the absence of a publicly available Persian medical language model, we opted to evaluate our model against general-purpose language models to establish a baseline for performance. This comparison allows us to assess the efficacy of our specialized model in handling medical-related queries in Persian. Importantly, all models used for comparison were selected based on their suitability for small, runnable environments on home devices, addressing privacy concerns prevalent in the medical domain. Additionally, we compared our model with a pipeline alternative in our evaluation, which consists of a series of processes: first, a translator converts the user’s query from Persian to English; next, this English query is input into an English medical language model; and finally, the response generated by the English model is translated back from English to Persian. By contrasting our model with both general-purpose language models and this pipeline alternative, we aim to demonstrate the advantages and specific capabilities of our small Persian medical language model in addressing the unique challenges of medical language processing within the Persian language context.  [^3] You can see the detailed result at https://github.com/Mehrdadghassabi/Gaokerena
@@ VI-A Comparison with general purpose language models
As you can see in Table III, our model achieved remarkable success by surpassing the passing score of the Iranian Basic Medical Sciences Entrance Exam, which stands at 36%, making it the first Persian language model with fewer than 8 billion parameters to pass this exam. Furthermore, our model demonstrated improvements on the translated MMLU dataset, achieving not only higher average scores but also excelling across most sub-categories, thereby showcasing its effectiveness in understanding and generating medical knowledge in the Persian language. Notably, the reason PersianMind has a low inference time, as shown in Table III, is that it tends to generate very short responses compared to other models, producing the end-of-sentence token much sooner. In addition to our multiple choice question answering evaluation, we also utilized GPT-4o [32] as an evaluator for free-form question answering. We provided the test set from the MF3QA dataset to both the competing language model and our model. As shown in Figure 3, GPT-4o predominantly preferred the responses generated by our model over those from the other three language models tested. This indicates that our model delivers high-quality responses, as judged by an advanced language model.
TABLE III: our model performance in comparison with general purpose language models
Gao- aya-
kerena expanse-8b Qwen2.5 PersianMind
@@
inference time $\approx 10s$ $\approx 8s$ $\approx 15s$ $\approx 2s$

### v3

@@ Abstract
The rapid advancement of language models has demonstrated the potential of artificial intelligence in the healthcare industry. However, small language models struggle with specialized domains in low-resource languages like Persian. While numerous medical-domain websites exist in Persian, no curated dataset or corpus has been available—making ours the first of its kind. This study explores the enhancement of medical knowledge in a small language model by leveraging accessible online data, including a crawled corpus from medical magazines and a dataset of real doctor-patient Q&A pairs.
We fine-tuned a baseline model using our curated data to improve its medical knowledge. Benchmark evaluations demonstrate that the fine-tuned model achieves improved accuracy in medical question answering and provides better responses compared to its baseline. Notably, the trained model successfully passed the Iranian Basic Medical Science Entrance Exam, taken in September 2023, and improved Persian-translated MMLU accuracy by an average of 2.67%. This work highlights the potential of leveraging open-access online data to enrich small language models in medical fields, providing a novel solution for Persian medical AI applications suitable for resource-constrained environments.
@@ I Introduction
In response to these challenges, we present a novel approach with our model, Gaokerena-V  [^1] Our language model is named after Gaokerena, an ancient Persian mythological tree believed to possess healing properties and grant immortality to those who consume its fruit. , which fine-tunes a baseline model, aya-expanse-8b [5], on a crawled dataset comprising a Persian medical corpus and a newly introduced free-form Farsi medical question-answering dataset. The crawled corpus contained 90 million tokens, of which 60% was utilized for training the model. Additionally, we introduced a new dataset named MF3QA, which contains 20,000 patient-doctor interaction pairs, further enhancing the model’s ability to handle diverse medical queries. aya-expanse-8b was specifically chosen as the baseline model due to its strong understanding of Persian language grammar, making it an ideal foundation for a Persian language model. Importantly, our model, corpus, and datasets are all open-source, promoting transparency and collaboration within the research community. This development aims to enhance access to Persian medical information and support secure, efficient interactions within the healthcare environment. By bridging the existing gaps in resources and leveraging advancements in NLP, our work contributes to the growing landscape of AI in medicine, particularly for Persian-speaking users.
@@ II-B Related Works in Persian
As previously mentioned, there has been limited research focused on Persian medical language models, highlighting a significant gap in resources for the Persian-speaking medical community. Furthermore, existing works on Persian medical question-answering systems are entirely closed-source regarding their datasets, models, and codebases. This lack of public resources leaves the field largely underexplored, presenting researchers with an almost blank slate to build upon. On the other hand, all of these efforts have primarily concentrated on extractive solutions, which aim to retrieve relevant information from predefined sources, rather than employing generative approaches capable of producing context-aware responses. In contrast to all previous studies in the Persian language domain, our approach is generative, not extractive, leveraging the internal knowledge of a language model to produce answers by synthesizing context and understanding. Generative question-answering systems rely on the internal knowledge of a language model, enabling them to produce answers by synthesizing context and understanding, whereas extractive question-answering systems retrieve answers from expert-generated data. For a detailed comparison of these two approaches, please refer to Table I.
TABLE I: Comparison of extractive question answering systems and generative question answering systems.
extractive QA systems generative QA systems
relies on information retrieval generative AI
using LMs encoder based LMs decoder based &
encoder-decoder based LMs
architecture expert generated data + a generative
an information language model
retrieval system
advantages reliability flexibility
disadvantages data dependency hallucination
Perhaps the most notable effort in the Persian medical language models field is Sina-BERT [4], which involved training a BERT model [11] using a crawled corpus alongside Persian annotated datasets specifically developed for various tasks, including medical question answering, medical sentiment analysis, and medical question retrieval. Sina-BERT is the most similar work to ours among Persian language focused efforts; however, it differs in that it uses a BERT model—an encoder-based language model—as its baseline. This choice limits its capability for generative AI tasks, as BERT is primarily designed for understanding and extracting information rather than generating answers.
@@ III Baseline Model
We have chosen aya-expanse-8b as our baseline model primarily due to the lack of open-source Persian medical language models, which necessitates the use of a general-purpose language model. While there are several multilingual options available, including aya-expanse [5], gemma-2 [15], qwen2.5 [16], and persianmind [17], we have determined that aya-expanse is the most suitable choice for our needs. One key reason is that the training data for the other models predominantly consists of non-Persian languages, leading to biases that may result in the generation of non-Persian characters, even when we explicitly instruct the model to use only Persian. In contrast, aya-expanse demonstrates a robust understanding of Persian grammar and produces grammatically rich Persian text, making it a more reliable option for our research. Additionally, we conducted an ablation study by fine-tuning qwen2.5-7b, and the results further validated our decision to select aya-expanse-8b as the baseline model. For more details, please refer to the ablation study section. Furthermore, if we merge our updated parameters into aya-vision [18], another model from the aya family, instead of aya-expanse, we gain the capability to incorporate medical images such as MRIs and CT scans as inputs, thereby enhancing our model’s ability to process complex medical questions.
@@ IV-A Corpus
I. Garcia Ferrero et al. [19] collected medical corpora dedicated to four languages (English, French, Spanish, and Italian), which can be compared to ours, as shown in Table II. The accompanying Figure 1 illustrates the share of each magazine within our corpus, effectively highlighting the diversity of sources and underscoring the need to address gaps in available resources to foster innovation and improve health-related applications.
TABLE II: Comparison of Our Corpus with Corpora Collected by I. Garcia Ferrero et al.
@@ IV-B Dataset
The collection of a real-world doctor-patient question-answering dataset is crucial for enhancing the capabilities of language models in the healthcare domain. Such a dataset allows models to learn valuable information derived from authentic interactions between healthcare providers and patients. By analyzing these real-world exchanges, language models can grasp the nuances of medical terminology, patient concerns, and the context surrounding healthcare inquiries. Furthermore, this dataset equips models with the ability to learn not just the factual content of responses but also the appropriate structure and tone for answering questions. This dual learning process is essential, as it enables the model to generate accurate, empathetic, and contextually relevant responses, ultimately improving patient communication and support in medical environments. In this context, Yang Liu [20] highlights several real-world doctor-patient question-answering datasets in his survey, a comparison of these datasets with ours can be found in Table III. In an era where technology increasingly aids healthcare, a robust doctor-patient dataset stands as a foundational element in training models that can effectively contribute to better healthcare delivery.
TABLE III: Comparison of Our dataset with others
@@
In our research, we crawled more than 180,000 question-answer pairs from Persian medical forums using the BeautifulSoup library 333you can see the crawling codes at https://github.com/Mehrdadghassabi/Gaokerena-V/tree/main/dataset/MF3QA/crawling_and_filtering , employing both manual and automatic filtering methods to refine the dataset through a laborious cleaning process. This approach is similar to the work done by Yunxiang Li et al. [6]. In their article on the Chat Doctor medical language model, where they extracted data from English medical forums, Yunxiang Li discarded about half of the question-answer pairs based solely on the length of the answers, as shorter responses are generally inadequate for training a model and can lead the model to learn to provide brief answers. However, we faced a greater challenge; Persian doctors tend to provide much shorter answers compared to their English counterparts, resulting in the necessity to discard over 80% of our question-answer records to ensure quality and relevance for our training purposes.
While Yunxiang Li only cleaned his question-answer pairs based on answer length, we used a more sophisticated process to do so. Firstly, we omitted answers with short lengths (less than 50 tokens) randomly in the training and development sets to shift the token-count histogram to the right. This random omission was not applied to the test set; instead, we meticulously reviewed each question-answer pair in the test set to ensure validity and accuracy. Additionally, we carefully read all question-answer pairs to filter out spam, duplicates, or low-quality records, ensuring that all remaining data contained high-quality information. We also removed any personal data to ensure compliance with privacy standards.
As you can see in the Figure 2 to create our dataset, we utilized patient-doctor interactions from the drhast and niniban platforms for the training split.It is important to note that drhast does not provide all of its doctor-patient interaction records on its site; it only offers access to the last 2,000 records. Additionally, each record is linked to 100 related records, complicating the crawling process. To address this challenge, we treated their data as a graph and performed a breadth-first search (BFS), which took about two weeks to extract 120,000 records out of a total of 200,000. The BFS we performed involved using a “for” loop where we selected one of the available 2,000 records as the root node and iteratively explored its linked records. By repeatedly performing BFS with new roots, we ensured broader coverage of the dataset. To prevent redundancy, we filtered out duplicated crawled records after each iteration. It is worth noting that the last 2,000 records on drhast are dynamic and change over time as new questions are asked, allowing us to use newly added records as roots for subsequent BFS runs. This iterative process, while time-consuming, ensured we could extract a substantial portion of the dataset within the two-week period. For the test set, we used the doctor-yab and isovisit sites, ensuring diversity by translating the K-QA question-answering dataset. [27] and appending it to our test split. This comprehensive approach not only enriched our dataset but also underscored the importance of real-world doctor-patient interactions in training effective language models.
@@ V-A Fine tuning
We fine-tuned the 8-billion-parameter checkpoint of the aya-expanse model on 60% of our corpus, focusing on minimizing resource usage. To ensure an efficient fine-tuning process, we employed gradient checkpointing and a small batch size of 2,reducing memory requirements during training. Additionally,we used gradient accumulation steps of 16, effectively increasing the overall batch size to 32 and enabling stable training dynamics. To further reduce the memory usage of our fine-tuning process, we leveraged Low Rank Adaptation (LoRA) [28] to significantly reduce the number of trainable parameters. Specifically, we implemented a rank of 8, an alpha value of 16, and a dropout rate of 5%, applying the LoRA weights to all trainable parameters of each transformer layer.To further optimize the process, we employed efficient tokenization and memory-aware training techniques. The tokenization process divided the input text into manageable token sequences, ensuring consistent input and label structures by truncating, padding, and handling overflowed tokens to maintain contextual integrity within a fixed context length. This streamlined preparation, coupled with the LoRA-based fine-tuning, was further enhanced by Flash Attention 2 [29].By minimizing memory overhead, Flash Attention 2 allowed us to handle longer context lengths and larger batch sizes efficiently, enabling effective fine-tuning for the next-token prediction objective while balancing computational efficiency with model performance. Figure 3 presents the loss curve during the fine-tuning stage. The curve demonstrates a smooth and consistent decrease in training loss, indicating stable convergence without signs of divergence or severe overfitting.
Figure 3: Training loss curve during the fine-tuning stage.
@@ V-B Instruction tuning
Following the fine-tuning stage, we performed instruction-tuning on the fine-tuned model using our crawled medical free-form Farsi question-answering (MF3QA) dataset.As you can see in table IV this stage utilized the default template of the aya-expanse model and retained the same techniques and almost the same hyperparameters as the fine-tuning stage, with a few adjustments. These adjustments were made to prevent the model from overfitting on the MF3QA dataset and to ensure it primarily learned the speaking style of a doctor. Specifically, we employed the LoRA method with a rank of 2, an alpha value of 2, and a dropout rate of 0.4, while increasing the weight decay to 0.5 (instead of 0.1). The instruction-tuning process was conducted for a single epoch, enabling the model to better understand and generate responses tailored to Farsi question-answering tasks. This targeted optimization further refined the model’s capabilities, enhancing its effectiveness on our specific dataset. Figure 5 illustrates the training loss during instruction-tuning. Due to the smaller dataset size and single-epoch training, the loss converges more quickly, reflecting the model’s adaptation to the task-specific speaking style without overfitting.
TABLE IV: Training hyperparameters
Fine-tuning Instruction-tuning
Number of epochs 1 1
Batch size 2 2
Number of gradient
accumulation steps 16 16
Optimizer AdamW AdamW
Learning rate 5E-4 5E-4
Maximum gradient norm 0.3 0.3
Warmup ratio 0.03 0.03
Weight decay rate 0.1 0.5
Maximum context length 1024 1024
Padding strategy left side padding left side padding
Lora rank 8 2
Lora alpha 16 2
Lora dropout rate 0.05 0.4
Target modules all linear layers all linear layers
Figure 4: Training loss curve during the instruction-tuning stage.
@@ V-C Carbon Footprint
The carbon footprint of our model optimization—including both fine-tuning and instruction-tuning—was estimated based on hardware specifications and operational duration. The process ran for a combined total of 19 hours on NVIDIA A100 PCIe 40GB GPUs hosted in Google Cloud Platform’s asia-east1 region, utilizing a Google Colab Pro+ account for the hardware resources. Assuming a typical power consumption of 250 watts per GPU, the total energy used was 4.75 KWh (250 watts × 19 hours). Using the carbon intensity factor of the asia-east1 grid (0.56 kilograms of CO2 equivalent per kWh), this translates to 2.66 kilograms of CO2 equivalent emitted during the tuning process [30].
@@ VI Results
In the absence of a publicly available Persian medical language model, we opted to evaluate our model against general-purpose language models to establish a baseline for performance. This comparison allows us to assess the efficacy of our specialized model in handling medical-related queries in Persian. Importantly, all models used for comparison were selected based on their suitability for small, runnable environments on home devices, addressing privacy concerns prevalent in the medical domain. Additionally, we compared our model with a pipeline alternative in our evaluation, which consists of a series of processes: first, a translator converts the user’s query from Persian to English; next, this English query is input into an English medical language model; and finally, the response generated by the English model is translated back from English to Persian. By contrasting our model with both general-purpose language models and this pipeline alternative, we aim to demonstrate the advantages and specific capabilities of our small Persian medical language model in addressing the unique challenges of medical language processing within the Persian language context.  [^4] You can see the detailed result at https://github.com/Mehrdadghassabi/Gaokerena-V
@@ VI-A Comparison with general purpose language models
As you can see in Table V, our model achieved remarkable success by surpassing the passing score of the Iranian Basic Medical Sciences Entrance Exam, which stands at 36%, making it the first Persian language model with fewer than 8 billion parameters to pass this exam. Furthermore, our model demonstrated improvements on the translated MMLU dataset, achieving not only higher average scores but also excelling across most sub-categories, thereby showcasing its effectiveness in understanding and generating medical knowledge in the Persian language. Notably, the reason PersianMind has a low inference time, as shown in Table V, is that it tends to generate very short responses compared to other models, producing the end-of-sentence token much sooner. In addition to our multiple-choice question-answering evaluation, we also utilized GPT-4o [32] as an evaluator for free-form question answering. We provided the test set from the MF3QA dataset (which includes 2,000 records) to both the competing language model and our model. As shown in Figure 5, GPT-4o predominantly preferred the responses generated by our model over those from the other three language models tested. This indicates that our model delivers high-quality responses, as judged by an advanced language model. Moreover, GPT-4o was prompted to prefer responses that provide more accurate medical information while avoiding incorrect medical information, further validating the reliability and precision of our model in generating medically accurate answers. 555you can see the prompt at https://github.com/Mehrdadghassabi/Gaokerena-V/blob/main/evaluation/free_form_qa/gaokerena_vs_aya/Untitled20.ipynb
TABLE V: our model performance in comparison with general purpose language models
gao aya-
kerena-V expanse-8b Qwen2.5 PersianMind
@@
inference time $\approx 10s$ $\approx 10s$ $\approx 15s$ $\approx 2s$